# RoCoDA: Counterfactual Data Augmentation for Data-Efficient Robot Learning from Demonstrations

### v1

@@ II Related Work
Behavior Cloning (BC) is a fundamental approach in imitation learning where a policy is trained to mimic expert demonstrations by directly mapping states to actions. Several recent BC methods have enabled robots to execute complex tasks by mimicking teleoperation data. [1, 10] classify 3D subgoals for an end-effector to achieve, and [2, 4, 5] predict latent actions using a CVAE, VQ-VAE, and causal convolution, respectively. These methods have proven extremely effective at fitting fine-grained, multi-modal action distributions in their training data.
@@ III A Unified Perspective on Data Generation
$f$ is $G$-equivariant if, $\forall$ $x\in X$ and $g\in G$, $f(g\cdot x)=g\cdot f(x)$, where $\cdot$ denotes the group action of $G$ on the respective spaces. Equivariance ensures that applying a group transformation before or after $f$ yields consistent results under the group action on $Y$.
@@ III-A Causality: Data Generation with Counterfactuals
$\pi\left(g\cdot\left(s_{C},s_{I}\right)\right)=\pi\left(s_{C}\right),\quad\forall g\in G$
@@ IV-A Counterfactual Data Augmentation
Notably, the state can be factored into independent subsets, the data can be augmented by sampling in the product space of individual subsets [21].
@@ V Results
RQ4: What emergent behaviors arise when policies are constrained to focus solely on task-relevant information?
@@ 3. Scaling Synthetic Data
4. Qualitative Analysis
To explore RQ4, we identify emergent behaviors that the policy was not explicitly trained to exhibit. One such behavior is re-grasping, where the policy recovers from a failed grasp attempt by reopening the gripper and attempting to grasp again (Figure 4). We observed in rollouts that once the gripper is fully closed, the robot fails to continue the task as this action is not in the training distribution (grippers are only closed partially when picking up objects). During sequences where the robot is in a transition phase and not interacting with other entities, local factorization of the gripper position and augmenting the states of the gripper positions can be performed. This learned behavior indicates a level of adaptability and error recovery not present in the training data.
@@ VI Conclusion
Appendix
VI-A Assumptions & Limitations
  1. 1.
    Generating demonstrations from human source. We use the same assumptions as [13] to generate new demonstrations.
  2. 2.
    Access to simulation states. We use information about simulation states to create new demonstrations and to create causally augmented data.
  3. 3.
    Subtask boundaries are determined by changes in gripper state. Defining the subtask boundaries by changes in the gripper state limits the the environments we experiment on. Environments that involve skills such as pushing or pulling objects would require using a different heuristic to divide subtasks.
  4. 4.
    Causal Relationships are known. The causal relationships between entities are assumed to be known by a domain expert. For each subtask, each timestep is labeled with its corresponding causal graph.
VI-B Additional Related Works
Data Augmentation
Data augmentation has played a critical role in the success of machine learning models by artificially expanding the size of training datasets and improving model generalization. Early work in this area applied basic transformations such as rotation, scaling, translation, and distortion to image datasets like MNIST [28], laying the groundwork for modern augmentation methods. The success of convolutional neural networks in image classification tasks further highlighted the importance of data augmentation. AlexNet [29] employed random cropping and horizontal flipping to improve image classification performance. Subsequent methods like Cutout [30] and Mixup [31] introduced techniques for robustness to occlusion and linear interpolation between samples, respectively.
Generative Data Synthesis
As generative models continue to gain popularity [32, 33], researchers have found multiple ways to use them to generate synthetic data for robotics. GenAug [18] Utilizes generative models like Stable Diffusion to alter textures and generate diverse visual samples, improving model robustness to appearance variations. ROSIE [19] Performs infilling using generative models to complete partially observed states and generate data with diverse backgrounds. SynthER [20] trains diffusion models on existing datasets to produce synthetic demonstrations, thus increasing the diversity and size of the training data. While some approaches [20, 34, 35, 36, 37] train diffusion models on existing datasets to produce synthetic demonstrations, increasing the diversity and size of the training data, these methods often focus on high-level semantic variations and may not adequately address low-level spurious correlations in state-action relationships. In this work, we establish a framework that leverages explicit causal and geometric augmentations. Exploring generative models as a complementary strategy is left for future work.
Image resolution 240x320x3
Learning rate 1e-4
Batch size 16
# of Encoder layers 4
# of Decoder layers 7
Feedforward dimensions 1600
Hidden dimension 256
# of Heads 8
Chunk size 15
Beta 10
Dropout 0.1
Table V: Hyperparameters of ACT
Algorithm 1 Offline Causal Augmentation
1:Dataset of demonstrations $\mathcal{D}=\{\tau_{i}\}$, where $\tau_{i}=\{s_{t},a_{t}\}_{t=1}^{T_{i}}$, causal graphs $\mathcal{G}$ for task.
2:Initialize $\mathcal{D}_{aug}\leftarrow\mathcal{D}$
3:for each trajectory $\tau\in\mathcal{D}$ do
4:     for each timestep $t\in[1,T_{i}]$ do
5:         Retrieve causal graph $\mathcal{G}$ for current phase
6:         Identify causally independent partitions
7:          of entities in $\mathcal{G}$
8:         for each partition do
9:              Sample a new partition from another
10:               trajectory in the same causal phase
11:              Replace the current partition with the
12:               new partition
13:              Render new observation
14:     Add augmented $\tau$ to $\mathcal{D}_{aug}$
15:return Augmented dataset $\mathcal{D}_{aug}$
Figure 5: We conduct experiments across five environments, each with a specific goal: (a) stack three blocks in a specified color order, (b) transport a hammer from the left bin to the top-right bin, (c) assemble a structure by connecting pieces in a particular sequence and orientation, (d) place a coffee pod into a coffee maker and close the lid, and (e) pick up a tomato sauce and place it into a basket.
VI-C Task Descriptions
In this section we provide more detail of tasks used in our experiments (Figure 5). For our experiments we use a panda arm for data collection and policy rollout. We use source data in collected in $D_{0}$ variations of Three Block Stack, Three Piece Assembly and Coffee.
Three Block Stack: The robot must stack three distinct colored blocks in order of green, red, and then the blue cube on top. We generate trajectories for the $D_{1}$ [13] variation of a larger spatial distribution of the blocks on the table.
Three Piece Assembly: This task requires precision and correct orientation, making it more challenging than simple block stacking due to the importance of part alignment. This task involves building a structure by assembling three pieces in a precise sequence and orientation. We generate trajectories for the $D_{2}$ [13] variation of the environment where all three pieces have a top-down rotation variation and all pieces can be initialized in different positions. We note that using MimicGen yielded a low success rate of generating trajectories. As MimicGen can generate biased data [13] low performance may be due to biased object configurations in the generated dataset.
Coffee Task: This task tests interaction with articulated objects and sequential dependencies, emphasizing the handling of non-rigid dynamics involving placing a coffee pod into a coffee machine and closing the lid. Different from the source demonstrations the $D_{2}$ [13] demonstrations have a larger region and rotational variation and the coffee pod and coffee machine initialized positions are switched.
Libero-Object: We use a task from the Libero-Object dataset [27] where the robot must pick up a specific object (e.g., tomato sauce) and place it in a basket. We modify this task during evaluation to measure generalization to unseen objects (distractors), textures, and positions. These experiments were trained with a subset of distractor objects. The evaluations on textures and positions were conducted with the same subset of distractors.
Transport: This task tests coordination between multiple agents and temporal sequencing. Transport [15] involves two robotic arms. The left arm removes a lid and picks up a hammer, while the right arm moves a cube from a back bin to a front bin, takes the hammer from the left arm, and places it in another bin. The hammer, cube, bins, and lid each vary slightly in position within a defined region, with the hammer and cube also having variations in rotation. For this task we examine the effect of causal augmentation versus standard augmentation on complex, long-horizon tasks.
VI-D Causal Relationships for Tasks
In our experiments, we use spatial information of objects as states to be augmented. In tasks such as block stacking, where the stacking order depends on block color, we could further factorize object states (e.g. color, texture, etc) to enhance augmentation. However, we simplify the causal structure by not factoring object states in this way. We specify causal relationships for each subtask and provide causal graphs for Three Block Stack and Coffee in Figure 7. Subtasks are defined based on changes in the robot’s gripper state as mentioned in Section III. When sampling data, we use the corresponding causal graph to sample new states. For instance, the third phase of Three Block Stack has two independent partitions: (1) Robot and Cube C and (2) Cube A and Cube B. As the states of Cube A and B are irrelevant to the robot’s movement to Cube C, the states of Cube A and B can be sampled from another trajectory of the same subtask phase without affecting the causal relationship. We additionally factorize the robot’s gripper position and augment the gripper position randomly while the robot transits to pick up objects.
Figure 6: Success Rate vs ratio of demonstration data to synthetic data. The addition of causal augmentation improves performance. As the amount of synthetic data is increased, performance gains diminish.
VI-E Training Details
Across all experiments, we keep training hyperparameters unchanged. We use hyperparameters similar to [2], but decrease the image resolution, feedforward dimensions, and hidden dimensions by a factor of two. We additionally double the batch size and use a chunk size of 15 and do not use temporal ensembling (Table V). We train Three Block Stack, Three Piece Assembly, and Coffee datasets using the agentview and robot0_eye_in_hand camera views. As for the Libero Object dataset, we train on the agentview camera view. While for the Transport we experiment with single agentview and multicamera view with robot0_eye_in_hand, robot1_eye_in_hand, shouldercamera0 and shouldercamera1 (Table II). Similar to [3], actions are positional control and use a 6D rotation representation.
Figure 7: Causal relationships for each subtask. (a) Three Block Stack has four subtasks: grasp cube A, place onto cube B, grasp cube C, place onto cubes A and B. Similarly Three Piece Assembly follows the same general causal relationship as Three Block Stack. (b) For Coffee, there are two subtasks: grasp the coffee pod, then insert coffee pod into the holder before closing the lid. The colors represent partitions of independent entities where entities of the same color are causally dependent and entities of different colors are causally independent.
VI-F Scaling Synthetic Dataset Size
We experiment with scaling the synthetic dataset size to understand the impact of data quantity on generalization. While data augmentation effectively increases the dataset and improves a model’s generalization, there are diminishing returns. Similar to [22] we explore increasing the dataset size and assess the ratio of demonstration data to causally augmented data. Separate from Section IV-A where data is sampled with some probability of applying causal augmentation, we augment offline as described in Algorithm 1.
The performance of scaling the dataset is evaluated on Three Block Stack and Coffee using the same hyperparameters. We produce 500 demonstrations for each environment using $SE(3)$ equivariant data augmentation, expand the data up to 10x through causal augmentation, and train using the number of gradient steps fixed at one million. We report our performance as the maximum success rate for all policy evaluations on 3 separate seeds as done in [15]. In the Three Block Stack environment, performance improves by approximately $20\%$ when adding synthetic data up to a ratio of 1:1 (Figure 6). However, as the dataset is expanded beyond the 1:1 ratio, the performance gains diminish with success rates falling within the margin of error. This indicates that adding more synthetic data beyond a 1:1 ratio does not result in substantial improvement and suggests some flexibility in the synthetic-to-real ratio. In certain tasks, we observed that counterfactual data augmentation sometimes led to out-of-distribution samples that reduced performance, especially at higher synthetic ratios (e.g., 1:5 in Coffee). This may be attributed to the model’s sensitivity to shifts in the distribution when synthetic data dominates. Similarly, this aligns with discussion of RQ3 in Section V-B as Three Block Stack benefits largely from augmentation while Coffee has marginal performance gains. We again attribute this to the complexity of each task – notice that there are a total of 8 causally independent state partitions throughout the execution of Three Block Stack, compared to only 2 for Coffee.

### v2

@@ II Related Work
Behavior Cloning (BC) is a fundamental approach in imitation learning where a policy is trained to mimic expert demonstrations by directly mapping states to actions. Several recent BC methods have enabled robots to execute complex tasks by mimicking teleoperation data. [1, 10] classify 3D subgoals for an end-effector to achieve, and [2, 11] predict latent actions using a CVAE, while [3, 4, 5] use diffusion, VQ-VAE, and causal convolution, respectively. These methods have proven extremely effective at fitting fine-grained, multi-modal action distributions in their training data.
@@ III A Unified Perspective on Data Generation
The function $f:X\rightarrow Y$ is $G$-equivariant if $\forall$ $x\in X$ and $g\in G$, $f(g\cdot x)=g\cdot f(x)$, where $\cdot$ denotes the group action of $G$ on the respective spaces. Equivariance ensures that applying a group transformation before or after $f$ yields consistent results under the group action on $Y$.
@@ III-A Causality: Data Generation with Counterfactuals
$\pi\left(g\cdot\left(s_{C},s_{I}\right)\right)=\pi\left(s_{C},g\cdot s_{I}\right)=\pi\left(s_{C}\right),\quad\forall g\in G$
@@ IV-A Counterfactual Data Augmentation
Notably, the state can be factored into independent subsets, the data can be augmented by sampling in the product space of individual subsets [22]. In the context of robotic manipulation, sampling from the product space of factorized state partitions involves selecting object states from different demonstrations while ensuring consistency with the joint distribution of dependent states.
@@ 3. Scaling Synthetic Data
V-C Emergent capabilities
We identify emergent behaviors that the policy was not explicitly trained to exhibit. One such behavior is re-grasping, where the policy recovers from a failed grasp attempt by reopening the gripper and attempting to grasp again (Figure 4). We observed in rollouts that once the gripper is fully closed, the robot fails to continue the task as this action is not in the training distribution (grippers are only closed partially when picking up objects). During sequences where the robot is in a transition phase and not interacting with other entities, local factorization of the gripper position and augmenting the states of the gripper positions can be performed. This learned behavior indicates a level of adaptability and error recovery not present in the training data.